OpenStudy (anonymous):

The perimeter of a rectangle is 38m. Find the dimensions of the rectangle that will contain the greatest area.

OpenStudy (anonymous):

can you help me?

OpenStudy (anonymous):

Let the sides of the rectangle be x and y 2x+2y=38 x+y=19 Area: xy = x(19-x) AM/GM inequality x(19-x) <= (x+19-x)²/4 = 19²/4 x=y=19/2 gives max area 361/4 and occurs when rectangle is a square

OpenStudy (anonymous):

another method.. let x be the width and y length of the rectangle P=2(x+y) from the given info p=38 2(x+y)=38 (divide both sides by 2) x+y=19 (subtract x from both sides) y=19-x area=x*y and we want x*y=max from the above calculation y=19-x x*y=x*(19-x)=max \[x(19-x)=19x-x ^{2}=\max\] by completing the square method \[19x-x ^{2}=-(x+\frac{ 19 }{ 2 })+\frac{ 19 }{ 2 }\] now draw the graph -(x+\frac{ 19 }{ 2 })+\frac{ 19 }{ 2 }\]  or simply the vertex gives the maximum values of x and y. \[x=\frac{ 19 }{ 2 } \] and \[y=\frac{ 19 }{ 2 } \]